What is Van’t Hoff’s factor for a compound which undergoes tetramerization in an
organic solvent?

Grade:upto college level

5 Answers

Sunil Kumar FP
9 years ago
In case of association i (van’t Hoff factor) is given by
i = 1 – (1- 1/n)α, where
n= number of molecules associates = 4 (here),
α = fraction of molecules associates

i = 1 – (1- ¼)α = 1 – ¾ α
if α be 1 (complete association), then

i= 1 – ¾ = ¼
akul
6 years ago
1 percent solution of KCL was found to freeze at -0.46 degree . calculate vant hoff factor and degree of dissociation of kf for water is 1.86kg.mol
Sachit
6 years ago
van`t hoff factor = Number of moles of particle after association / Number of moles of particles before associationTherefore : 4X ➡ X4 and hence van`t hoff factor = 1/4
Kushagra Madhukar
3 years ago
Hello student
In case of association i (van’t Hoff factor) is given by
i = 1 – (1- 1/n)α
where n= number of molecules associates = 4 (here)
α = fraction of molecules associates
i = 1 – (1- ¼)α = 1 – ¾ α
if α be 1 (complete association), then i= 1 – ¾ = ¼

Hope it helps
Regards,
Kushagra
Yash Chourasiya
3 years ago
Dear Student

In the case of association i (van’t Hoff factor) is given by
i = 1 – (1- 1/n)α
where n= number of molecules associates = 4 (Because of tetramerization)
α = fraction of molecules associates
i = 1 – (1- ¼)α = 1 – ¾ α
if α be 1 (complete association),
then i= 1 – ¾ =¼

I hope this answer will help you.
Thanks & Regards
Yash Chourasiya